OpenStudy (anonymous):

Enter in standard form the equation of the line passing through the given point and having the given slope. B(2, - 5), m = -2/3

OpenStudy (saifoo.khan):

Use the short formula.. y - y1 = m(x-x1) y - (-5) = -2/3 (x - 2)

OpenStudy (saifoo.khan):

Solve.

OpenStudy (anonymous):

y+5 = 2/3x - 4/3

OpenStudy (saifoo.khan):

Correct.

OpenStudy (saifoo.khan):

Now bring that into standard form.

OpenStudy (anonymous):

2/3 x - y = i am stuck

OpenStudy (saifoo.khan):

make y the subject first.

OpenStudy (saifoo.khan):

y = 2/3 x -4/3 - 5

OpenStudy (anonymous):

y = -2/3 - 5

OpenStudy (saifoo.khan):

wrong.

OpenStudy (anonymous):

not sure what answer is then

OpenStudy (saifoo.khan):

Can you solve from here?